proof, sin t x csc t (tan t x cot t) = sec^2 t + csc^2 t
Answer by Alan3354(21583)

sin t x csc t (tan t x cot t) = sec^2 t + csc^2 t
----------
sin(t)*csc(t)*(tan(t)*cot(t)) = sec^2(t) + csc^(t)
sin*csc = 1
tan*cot = 1
----------
sec^2 + csc^2 = 1
1/cos^2 + 1/sin^2 = 1
(sin^2 + cos^2)/sin^2cos^2 = 1
sin^2 + cos^2 = 1
1/sin^2cos^2 = 1
sin^2cos^2 <> 1
It does not.